the age of ram's mother is 4 times the age of ram .five years ago,the sum of their ages was 30 years.find their present ages.
Answer by alicealc(293)

Ram's age = R
Ram's mother's age = M

The age of ram's mother is 4 times the age of ram:
M = 4R
Five years ago,the sum of their ages was 30 years:
M - 5 + R - 5 = 30
4R - 5 + R - 5 = 30
5R - 10 = 30
5R = 30 + 10
5R = 40
R = 40/5 = 8
M = 4R = 4*8= 32

so, at the present, Ram is 8 years old and Ram's mother is 32 years old